OpenStudy (unklerhaukus):

\[2y^{\prime\prime}+5y^\prime+2y=x^2\]

OpenStudy (unklerhaukus):

\[2y_c^{\prime\prime}+5y_c^\prime+2y_c=0\] \[2m^2+5m+2=0\]\[\Delta=5^2-4(2)(2)=9\]\[(m+2)(2m+1)=0\]\[m_{1,2}=-2,-\frac12\]\[y_c=Ae^{-2x}+Be^{-\frac x2}\] \[2y_p^{\prime\prime}+5y_p^\prime+2y_p=x^2\] \[y_p=Cx^2+Dx+E;\quad y_p^\prime=2Cx+D;\qquad y_p^{\prime\prime}=2C\]\[2(2C)+5(2Cx+D)+2(Cx^2+Dx+E)=x^2\]\[(2C)x^2+(10C+2D)x+(4C+5D+2E)=x^2\]\[2C=1:\qquad 10C+2D=0;\qquad 4C+5D+2E=0\]\[C=\frac12 ;\qquad D=-\frac52;\qquad 4\frac12-5\frac52=-2E\]\[-2E=-\frac{21}2;\qquad E=5\frac{1}4\]\[y_p=\frac12{x^2}-\frac52x+5\frac{1}4\] \[y=y_c+y_p\] \[y(x)=Ae^{-2x}+Be^{-\frac x2}+\frac12{x^2}-\frac52x+5\frac{1}4\]

OpenStudy (unklerhaukus):

ops sorry ive but up the wrong question, i think i have done this all right.

OpenStudy (unklerhaukus):

The is the question i ment to ask \[y^{\prime\prime}+8y^\prime+16y=25e^x\]

OpenStudy (unklerhaukus):

\[y_c^{\prime\prime}+8y_c^\prime+16y_c=0\] \[m^2+8m+16=0\]\[\Delta=(8)^2-4(1)(16)=0\]\[(m+4)^2=0\]\[m=-4\] \[y_c=(Ax+B)e^{-4x}\]

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}+8y_p^\prime+16y_p=25e^x\] \[y_p=Ae^{rx};\qquad y_p^\prime=rAe^{rx};\qquad y_p^{\prime\prime}=r^2Ae^{rx}\]\[(r^2Ae^{rx})+8(rAe^{rx})+16(Ae^{rx})=25e^x\]\[\left(r^2+8r+16\right)Ae^{rx}=25e^x\]\[\left(r^2+8r+16\right)Ae^{r}=25\]\[16Ae^r=25\]\[r=1;\qquad A=1\] \[y_p=e^{x}\] \[y(x)=(Ax+B)e^{-4x}+e^x\]

OpenStudy (unklerhaukus):

im not really sure what i did to get the constants , it kinda makes sense now i have them but \[\left(r^2+8r+16\right)Ae^{r}=25\] how do we find A,r

OpenStudy (unklerhaukus):

wouldn't A= 25/16 when r=0, or something

OpenStudy (unklerhaukus):

equating co-efficients or some such

OpenStudy (anonymous):

How can you cancel this: \[e^{rx} = e^x\]

OpenStudy (unklerhaukus):

i was suing an \(r\) for \(\alpha\)

OpenStudy (unklerhaukus):

havent i just timesed by \(e^{-x}\) @waterineyes

OpenStudy (unklerhaukus):

oh i see the mistake now

OpenStudy (anonymous):

take a look http://www.efunda.com/math/ode/linearode_undeterminedcoeff.cfm

OpenStudy (anonymous):

What you said I did not catch that..

OpenStudy (unklerhaukus):

from here\[\left(r^2+8r+16\right)Ae^{rx}=25e^x\]i should have \[\left(r^2+8r+16\right)Ae^{(r-1)x}=25\]

OpenStudy (unklerhaukus):

\[A=25\frac{e^{(1-r)x}}{(r-4)^2}\]

OpenStudy (unklerhaukus):

mm confused

OpenStudy (anonymous):

Now you are right...

OpenStudy (unklerhaukus):

what is the next step?

OpenStudy (anonymous):

actually the particular solution for 25e^x is Ae^x\[y_p=Ae^x\] \[y_p'=Ae^x\] \[y_p''=Ae^x\] \[25Ae^x = 25e^x\] \[A=1\] \[y=c_1e^{-4x}+c_2xe^{-4x}+e^x\]

OpenStudy (anonymous):

If I assume the particular solution as: \[y_p = ae^x\] \[y_p' = ae^x\] \[y_p'' = ae^x\] Substitute this: \[ae^x + 8ae^x + 16ae^x = 25e^x\] \[25ae^x = 25e^x\] So, a = 1.. \[y_p = e^x\] \[y(x)=(Ax+B)e^{−4x}+e^x\]

OpenStudy (unklerhaukus):

how come you can assume \(Ae^{x}\) as apposed to \(Ae^{rx}\)

OpenStudy (anonymous):

See your right hand side..

OpenStudy (anonymous):

We assume the particular solution according to the right hand side given so that we can equate and find the value of constants..

OpenStudy (unklerhaukus):

ah r=1 because it has to equal e^{1x}

OpenStudy (anonymous):

yes...

OpenStudy (unklerhaukus):

that is making some sense

OpenStudy (unklerhaukus):

in mukushla's table \(\gamma=1\)

OpenStudy (anonymous):

Always choose the particular solution in a way which will make it easy to equate coefficients..

OpenStudy (unklerhaukus):

i get it now thanks everyone

OpenStudy (anonymous):

Welcome dear..

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}+8y_p^\prime+16y_p=25e^x\] \[y_p=Ae^{x};\qquad y_p^\prime=Ae^{x};\qquad y_p^{\prime\prime}=Ae^{x}\]\[(Ae^{x})+8(Ae^{x})+16(Ae^{x})=25e^x\]\[\left(1+8+16\right)Ae^{x}=25e^x\]\[25A=25\]\[A=1\] \[y(x)=(Ax+B)e^{-4x}+e^x\]

OpenStudy (anonymous):

Yes this one is very true..

OpenStudy (unklerhaukus):

\[\color{red}\checkmark\]

OpenStudy (anonymous):

\[\Huge \color{green} { {\cancel{\heartsuit}}}\]

OpenStudy (unklerhaukus):

what does \[ \color{green} { {\cancel{\heartsuit}}}\]mean

OpenStudy (anonymous):

I want to make arrow in that heart..

OpenStudy (anonymous):

It is simply a sign of smile and happiness...

OpenStudy (anonymous):

how do you makes those symbols o.o also how to write fractions \[a/b\]

OpenStudy (unklerhaukus):

\frac{a}{b} \]

OpenStudy (unklerhaukus):

http://en.wikibooks.org/wiki/LaTeX/Mathematics

OpenStudy (anonymous):

O.o this is new ty vm